I honestly have NO idea how I am supposed to work this information

Let's shoot a laser, at a reflector on the moon. What diameter would the reflector have to be to capture the entire beam? distance to moon = 250,000 miles = 400,000 km = 4x10^10 cm.
Answer #1

Angular size = Diameter of reflector / Distance to moon

heta=rac{D}{X}

D= heta X

Beam angle rad 180

Distance from earth to moon X=4*10°km = 4 * 108 m

8 *-* 4 * 108-558 * 107 m = 5.58 * 104 km 180

Diameter of reflector must be D = 5.58 * 10 m = 5.58 * 104 km